Draw the net electric force on a +1-C charge placed at the center of: 

a. A line segment of length 1.0 m with a +1-C charge and a –1-C charge at its endpoints; 

b. An equilateral triangle of side length 1.0 m with alternating +1-C, –1-C, and +1-C charges 

placed at its vertices;

c. A square of side length 1.0 m with alternating +1-C, –1-C, +1-C, and –1-C charges placed 

at its vertices.

d. Calculate the magnitude and determine the direction of the net electric force on the +1-

C charge placed at the center of configurations in a, b, and c.


1
Expert's answer
2021-06-14T15:04:44-0400

Gives

Part

Mid point P(+1C)

Point A(+1C) and B(-1C )

Electric force at centre

"F_{Ap}=\frac{kq_1q_2}{r^2}"

Put value

"F_{AP}=\frac{9\times10^9\times1\times1}{0.5^2}=36\times10^9N"

"F_{BP}=-\frac{9\times10^9\times1\times1}{0.5^2}=-36\times10^9N"

"F_{net}=36\times10^9-36\times10^9=0N"

Part (b)

PointA (+1C) point B(-1C) pointC(+1c)

"F_{AP}=\frac{kq_1q_2}{r^2_{AP}}"

Put value

"F_{AP}=\frac{9\times10^9\times1\times1}{(\frac{1}{\sqrt3})^2}=27N"

"F_{BP}=-\frac{9\times10^9\times1\times1}{(\frac{1}{\sqrt3})^2}=-27N"

"F_{CP}=\frac{9\times10^9\times1\times1}{(\frac{1}{\sqrt3})^2}=27N"

"F_{net}=F_{AP}+F_{BP}+F_{CP}"

"F_{net}=27-27+27=27N"

Part(c)

Centre of force

"r_{AP}=r_{BP}=r_{CP}=r_{DP}=\frac{1}{\sqrt2}"

"F_{AP}=\frac{KQ_1Q_2}{r^2_Ap}"

Put value

"F_{AP}=\frac{9\times10^9\times1\times1}{(\frac{1}{\sqrt2})^2}=18N"

"F_{BP}=-\frac{9\times10^9\times1\times1}{(\frac{1}{\sqrt2})^2}=-18N"

"F_{CP}=\frac{9\times10^9\times1\times1}{(\frac{1}{\sqrt2})^2}=18N"

"F_{DP}=-\frac{9\times10^9\times1\times1}{(\frac{1}{\sqrt2})^2}=-18N"


"F_{net}=F_{AP}+F_{BP}+F_{CP}+F_{DP}"

Put value

"F_{net}=0N"

Part(d)

(a)

Due to centre for Line

Fnet=0N

Due to centre of triangle

Fnet=27N

Due to square at centre

Fnet =0N